# milliFlow: Scene Flow Estimation on mmWave Radar Point Cloud for Human Motion Sensing

### 12 Jan 2024 v4

@@ 4.2. Problem Formulation
Figure 2. mmWave-based scene flow network architecture.
4.3. Scene Flow Network
Following the state-of-the-arts (Ding et al., 2023; Wei et al., 2021; Kittenplon et al., 2021; Gojcic et al., 2021), we rely on end-to-end trainable deep neural networks to learn scene flow estimation offline and apply the trained model to infer scene flow for upcoming radar point clouds online. The overall scene flow network architecture can be seen in Fig. 2. In the following, we will first introduce five modules of our network in detail.
4.3.1. Local Feature Abstraction.
In our network, we adopt the set abstraction layer proposed in PointNet++ (Qi et al., 2017) as the basic learning layer to extract local point features. The working process of the local feature abstraction module can be found in Fig. 2. With two mmWave radar point clouds $\mathcal{X}$ and $\mathcal{Y}$ as input, we first encode their local features respectively using four parallel set abstraction (SA) layers (Qi et al., 2017). These SA layers have different grouping radii and each of them extracts the local feature $l_{i,s}$ for each radar point at a specific scale $s$. Such multi-scale local feature abstraction is more insensitive to the non-uniform density of radar point clouds. We further use a shared-weighted MLP to map the local feature $l_{i,s}$ into a higher-level representation $k_{i,s}=\mathrm{MLP}(l_{i,s};\theta_{s})$, where $\theta_{s}$ is the layer parameters for the scale $s$. For each point cloud, we concatenate the high-level local features from different scales and obtain the multi-scale local features $\mathcal{K}^{\mathcal{X}}$, $\mathcal{K}^{\mathcal{Y}}$ for two input point clouds. Besides, we also utilize a context feature extractor to encode the context features $\mathcal{K}^{\mathcal{X}}_{c}$ of point cloud $\mathcal{X}$. Note that this context extractor does not share weights with the local feature encoder but they have the same structure and hyperparameters. The extracted context feature is utilized in Sec. 4.3.3 to provide the context information to the flow embedding.
4.3.2. Global Feature Aggregation
As the first step, the local feature vector of each radar point is mapped into a scalar $w_{i}=\mathrm{MLP}(k_{i};\theta_{a})$, where $\theta_{a}$ is the learnable parameters for the attention mapping. To ensure numerical stability, we normalize all attention weights to sum to 1. Then, the global feature vector $g$ can be obtained by aggregating the local features of all points with a weighted sum operation, which is written as:
(3) $g=\sum_{i=1}^{N}w_{i}\cdot k_{i}$
Compared with the max-pooling operation that discards almost all information in local features, this attention-based aggregation can learn to dynamically adjust the weights and combine the information from all points. Finally, we concatenate this global vector to the local feature of each point and obtain the local-global representations $\mathcal{Z}^{\mathcal{X}},\mathcal{Z}^{\mathcal{Y}}$ for radar frames $\mathcal{X}$ and $\mathcal{Y}$. Similar to local feature abstraction, we also extract the global feature for the context feature $\mathcal{L}_{C}^{\mathcal{X}}$ using another MLP and denotes its output as $\mathcal{Z}^{\mathcal{X}}_{C}$.
4.3.3. Flow Embedding Generation
@@
To encode point motion between two radar frames, we leverage the point cloud cost volume (CV) layer in (Wu et al., 2020) to correlate features. Given the local-global features of two radar point clouds, i.e., $\mathcal{Z}^{\mathcal{X}},\mathcal{Z}^{\mathcal{Y}}$, we use the CV layer (Wu et al., 2020) to compute the correlation between them, as seen in Fig. 2. By aggregating the spatial relationship and feature similarities between two frames, the point motions are encoded into the cost volumes, denoted as $\mathcal{O}=\{o_{i}\}_{i=1}^{N}$. Thanks to the global feature aggregation, the holistic frame information can also be correlated, which yields more robust and stable costs. The CV layer enables us to decode the scene flow vector for each point as it contains the motion information between two frames. To further mix it with the context, we then stack the cost volumes $\mathcal{O}$, context features $\mathcal{L}_{C}^{\mathcal{X}}$ and pass the features into another local encoder to obtain the flow embedding =$\mathcal{B}=\{b_{i}\}_{i=1}^{N}$.
4.3.4. Temporal Information Propagation
@@
Directly propagating 2-dimension flow embedding from previous frames suffers from a) high computation overload, b) the change in the number of points between frames. To overcome these issues, we propose to temporally update its global vector instead of the flow embedding itself, as shown in Fig. 2. More specifically, we first follow the same process as in global feature abstraction to generate the global vector $g^{\mathcal{B}}$ of the flow embedding $\mathcal{B}$. Then we apply a GRU network (Cho et al., 2014) to update it as hidden states temporally and obtain the final global representation $h^{t}=\mathrm{GRU}(h^{t-1},g^{\mathcal{B}};\theta_{g})$, where $h^{t-1}$ is the global representation from the last frame and $\theta_{g}$ is the GRU network parameters. Lastly, we concatenate this updated global representation to each point in the flow embedding and denote the final features as $\mathcal{A}=\{a_{i}\}_{i=1}^{N}$.
4.3.5. Constrained Scene Flow Regression
@@
Given the final features $\mathcal{A}$ produced above, we can use an MLP-based flow regressor to decode per-point scene flow =$\mathcal{{F}}=\{{f}_{i}\}_{i=1}^{N}$. However, estimating unconstrained scene flow may lead to non-viable results, e.g., the magnitude of the flow vector exceeds the normal scale of human body movement. To constrain our predictions, we propose to clamp the estimated scene flow before returning it. We set a fixed threshold $\epsilon$ and constrain the scene flow on each axis to be within the range of [$-\epsilon,\epsilon$].
Figure 3. Automatic scene flow labelling pipeline.
4.4. Automatic Scene Flow Labelling
To cope with the human sensing scenarios where non-rigid human motion dominates, we propose a new cross-modal automatic scene flow labelling scheme, as shown in Fig. 3. With the help of the co-located RGB-D camera, we first label 3D human skeletons and then generate noisy pseudo scene flow labels with respect to the skeleton-based rigid-motion assumption.
@@ 4.4. Automatic Scene Flow Labelling
Human Skeleton Annotation. In human sensing applications, the human skeleton is usually characterised by the 3D position of its two endpoints, aka. keypoints. To quickly and conveniently annotate such keypoints, we refer to the RGB-D images recorded by the co-located depth camera in this work. Specifically, we first utilize an open-source pose estimation library (e.g. OpenPose (Cao et al., 2019)) to label 2D keypoints on RGB images and then uplift each 2D keypoint to 3D using its corresponding depth value. Then the human skeleton labels can be obtained using intrinsically connected 3D keypoints, as shown in Fig. 3.
@@
Keypoint-Based Label Filtering. Due to the errors induced by the keypoint estimator, our generated pseudo scene flow labels are prone to have much noise. To mitigate the impact of these errors, we propose to filter our labels based on the confidence scores and the displacement of estimated keypoints. First, we retrieve the confidence scores from the 2D keypoints estimator and filter out those keypoints whose confidence is lower than 0.5. Then, we calculate the 3D displacement for the left keypoints between two frames and further discard keypoints whose displacement is bigger than 0.5m. After the above two-stage filtering, we obtain a set of valid keypoints for each frame. Skeletons whose two keypoints are both valid are identified as valid and we generate a mask =$\mathcal{M}=\{m_{i}\in\{0,1\}\}_{i=1}^{N}$ to discard the pseudo scene flow labels associated with the invalid skeletons, where valid labels are marked as 1. This mask can be used to mitigate the impact of noises in our loss (c.f. Sec. 4.5).
4.5. Loss Function
@@
To help our network escape from the local minima, we propose to weigh points with large-scale movements more when computing the training loss. Specifically, we separately calculate the average loss for points whose scene flow label is larger than a threshold $\zeta$ and combine it with the average loss of other points. Our loss function can be written as:
(4) $\mathcal{L}=\alpha_{l}\mathcal{L}_{large}+\alpha_{s}\mathcal{L}_{small}$
Here, $\alpha_{l}$ and $\alpha_{s}$ are the hyperparameters to balance the impact of large- and small-scale movements. The error between the scene flow label and prediction is measured as the $L_{2}$ distance. By giving points with large-scale movement more weight in our loss, we can dilute the domination of small-scale points that constitute the majority, especially the static ones, to avoid the local minima.
5. Implementation
5.1. Dataset Collection
5.1.1. Sensing Platform.
As exhibited in Fig. 4 (a), we use a commercial Vayyar vTrigB imaging mmWave radar (vay, 2023) and a RealSense D455 depth camera (d45, 2023) to capture mmWave radar point clouds and RGB-D images respectively. The mmWave radar device is portable with a size of 10.4 cm $\times$ 8.5cm and a weight of 110g in total, emitting SFCW to the surroundings to receive the RF signals. With the default sensor setting, the range resolution of our mmWave radar is 9.35 cm, and the maximum range is 14 m, while the angular resolution is approximately 6.7 degrees. The internal data processing pipeline of this device follows the steps we illustrate in Sec. 2. The final output is a set of 3D points with per-point intensity values. For the depth camera, the RGB-D image size is set as 640$\times$480 and its depth measurement range is from 0.6m to 6m. Both sensors are fixed on a spoken collection board to ensure their relative position is unchanged after the initial calibration. We further mount this board on a tripod with a height of about 1.3m, as seen in Fig. 4 (a). During collection, we alternately query the data frame from two sensors and stored them in a PC, resulting in the synchronized mmWave radar point clouds and RGB-D images. The average frame rate over all collected data is $\sim$13.2Hz.
5.1.2. Collection Subject.
@@ 5.1. Dataset Collection
12 participants of various genders, ages, heights and weights are recruited for data collection in this work [^1] The study has received the ethical approval from XXX (no name for blind review), and participant consent forms were signed before the collection.. To evaluate the effectiveness of methods in different scenes, we choose three sites, as shown in Fig. 4 (c), including the hallway, parking lot and square as our human sensing experiment scenes. For each scene, each participant is asked to perform 5 ’in-set’ activities, namely bicep curl, arm swing, bowing, leg swing, and arm & leg swing, as seen in Fig. 4 (d), each of which lasts for about 20s. The collection platform is placed about $2.5-3.5$m away from the subject. Every subject is asked to wear a face mask to protect their identities from being recognised. After data collection, we manually crop each sequence to 200 frames for uniform activity distribution. As a result, the whole main dataset consists of $12\times 3\times 5\times 200=$ 36k frames in total. To test the generalization ability to unseen subjects, we divide the dataset into three parts by subject. 6 people are used for training, while 2 people are used as the validation set and the remaining 4 people are used for testing. Besides the main dataset, we also collect an ’out-of-set’ testing set by letting all people perform three ’out-of-set’ activities, including squatting, sitting and head bobbing as shown in Fig. 4 (d), in the hallway site to evaluate the generalization to unseen activities. This extra testing set is composed of $12\times 3\times 200=$ 7.2k frames.
@@
Figure 4. Collection setup, test environment, subject activities and an example of pseudo pose labels.
5.2. Evaluation Setup
5.2.1. Data Preprocessing.
5.2.2. Labelling.
We use our proposed automatic labelling scheme (c.f. Sec. 4.4) to annotate scene flow for training and validation samples. Specifically, we keep 14 keypoints from the off-the-shelf pose estimator to represent the subject pose following  (Xue et al., 2021; Ren et al., 2022), which results in 13 connections between intrinsically connected keypoints as seen in Fig. 4 (b). We then omit invalid skeletons with our two-stage label filtering and annotate the pseudo scene flow labels for points belonging to the valid ones. For our testing samples, we follow the same labelling pipeline but annotate all points. To make the testing labels accurate enough, we added manual inspection and correction for the estimated 2D keypoints.
For the human action recognition (HAR) task, we manually label the ground truth of human activity for each sequence during recording. To annotate human parsing (HP) labels, we utilize the point-skeleton affiliation information generated during scene flow labelling and label each point with the body part it belongs to. Specifically, as the skeleton labels generated during training are noisy, we also introduce the valid mask $\mathcal{M}$ (c.f. Section 4.4) to our HP task to discard invalid points for training. The testing labels for the human body part tracking (HBPT) task are ready to use as we already obtain the 3D keypoints during scene flow labelling.
5.2.3. Evaluation Metric.
We use the following evaluation metrics to quantify the performance on scene flow estimation and downstream human sensing tasks.
  • EPE3D (m). It computes the average 3D endpoint error $||f_{i}-f_{gt}||_{2}$ over all points in a frame. Following (Baur et al., 2021; Ding et al., 2022), we also report the errors separately for the moving and static points. Points with a ground truth flow vector larger than 0.01m are labelled as moving.
  • Acc3D. It measures the ratio of points in one frame that meets the relax/strict accuracy requirements, i.e., either EPE3D ¡ 0.025/0.05m or relative error ¡5%/10%.
5.2.4. Network and Training Details.
For the scene flow network, the grouping radii for four SA layers used for local or context feature extraction (c.f. Sec. 4.3.1) are [0.05, 0.1, 0.2, 0.4] meters. The numbers of local samples for them are set as [4, 8, 16, 32] and the dimension of the MLP used in and after each SA layer is [32, 32, 64], [64, 64, 64] respectively. In all global feature aggregation (c.f. Sec. 4.3.2), the MLP used for attention mapping is two-layer with a hidden dimension of 128. In the CV layer (c.f. Sec. 4.3.3), the number of neighbours for patch-to-patch aggregation is set as 8 and each MLP used to learn aggregation weights has three layers with hidden dimensions of [8, 8]. The local encoder used to generate flow embedding has the same hyperparameters as the former one except for its MLP which has the dimension of [512, 256, 64] in each SA layer. The dimension of the flow regressor MLP (c.f. Sec. 4.3.5) is [256, 128, 64, 3] and the threshold $\epsilon$ to constrain the output is set as $0.1m$.
To implement our temporal propagation (c.f. Sec. 4.3.4), we divide the long sequence of scene flow samples into many mini-clips with a length of 5 frames and train using mini-batches of them after shuffling. During inference, we re-initialize the hidden state of the GRU to zero vectors after 5 frames to fit the training pattern. The threshold $\zeta$ used for our loss function (c.f. Eq. 4.5) is fixed as 0.1m for all experiments while the weight hyperparameters $\alpha_{l}$ and $\alpha_{s}$ are set as 2 and 1. For all training in our experiments, we consistently use the Adam optimizer with an initial learning rate of 1e-3, which decays by 0.9 after each epoch. The performance on the validation set is used to determine when to stop the training process and to select the best model for each training experiment. More details on how to implement our downstream tasks can be found below in Sec. 7.
6. Scene Flow Evaluation
Figure 5. Qualitative scene flow estimation results. To make our results more readable, we project radar points and scene flow vectors onto the image plane for visualization. We use red for the ground truth while blue for ours.
6.1. Overall Performance
6.1.1. Compared to the state of the arts.
@@ 6.1. Overall Performance
FlowNet3D (Liu et al., 2019) 0.293 0.290 0.259 0.016 0.095
PPWC-Net (Wu et al., 2020) 0.171 0.181 0.128 0.138 0.179
Graph Prior (Pontes et al., 2020) 0.315 0.322 0.283 0.007 0.011
FLOT (Puy et al., 2020) 0.299 0.307 0.265 0.015 0.094
FlowStep3D (Kittenplon et al., 2021) 0.243 0.251 0.216 0.062 0.109
PV-RAFT (Wei et al., 2021) 0.161 0.170 0.107 0.179 0.292
RaFlow (Ding et al., 2022) 0.107 0.115 0.094 0.271 0.427
Bi-PFNet (Cheng and Ko, 2022) 0.159 0.168 0.111 0.153 0.264
@@
Table 1. Comparison of scene flow estimation results.
We first compare our network with state-of-the-art methods of point-based scene flow estimation. Our baselines include 7 deep learning-based methods (Liu et al., 2019; Wu et al., 2020; Kittenplon et al., 2021; Puy et al., 2020; Wei et al., 2021; Cheng and Ko, 2022; Ding et al., 2022) whose networks are also supervised with pseudo scene flow labels, and one non-learning-based method Graph Prior (Pontes et al., 2020) that solves scene flow via online optimization. The evaluation results on our testing set can be seen in Table 1. The scene flow network of milliFlow achieves a cm-level average EPE3D (i.e., 4.6cm) and a high relax Acc3D of 70.3%, ranking 1st on all metrics with a large margin compared to the baselines. This satisfactory performance proves the efficacy of our method to address the challenges existing in our scene flow task. Note that our network is trained with pseudo labels automatically generated using RGB-D images (c.f. Sec 4.4), which does not demand any manual annotation efforts.
6.1.2. Qualitative results.
Some examples of our scene flow outputs are demonstrated in Fig. 5. It can be seen that our method can produce reliable scene flow estimation for various activities performed by different subjects in three test environments. Despite the sparsity and noise of radar point clouds, our network can still effectively learn representative features thanks to our global feature aggregation (c.f. Sec. 4.3.2) and robustly regress scene flow vectors. As exhibited in our examples, some body parts are missing in radar point clouds, which impedes the scene flow inference process. We compensate for the temporal information shortage by utilizing the historical information to benefit the current scene flow estimation (c.f. Sec. 4.3.4), which yields sound testing performance across frames.
6.1.3. Runtime efficiency.
We test the runtime efficiency of our scene flow network on a single NVIDIA RTX 3090 GPU. Given sequential testing radar point clouds, we feed them one by one into our trained model for inference. As a result, our model has real-time performance with one inference step in 74ms ($\sim$ 13.5Hz). Moreover, the maximum allocated GPU memory is only 134 MB during inference. Such a low GPU memory consumption facilitates the parallel operation of scene flow and downstream networks.
6.2. Ablation Study
@@
Table 2. Breakdown results of our scene flow network.
To validate the effectiveness of some key components in our scene flow network, we disable them one by one and investigate the impact on the results. Specifically, four components are ablated from our network, including temporal propagation (TP), global aggregation (GA), context feature (CF) and constrained regression (CR). The ablation experiments results are exhibited in Table 2. Overall, the full version of our network (row (a)) yields the best results and each component helps to elevate the performance on each metric (from row (e) to (a)). First, our temporal propagation can effectively harness the information from previous frames to help scene flow estimation (c.f. Sec. 4.3.4), yielding a relatively 13.2% lower EPE3D and a 6.3% increase on strict Acc3D (comparing row (a) and (b)). Moreover, by leveraging our attention-based global feature aggregation (c.f. Sec. 4.3.2), the vulnerable local features can be largely augmented with holistic information, which brings another remarkable improvement over each metric (e.g. 13.1% on EPE3D). Besides, incorporating the context features into our flow embedding can help to preserve the context from the source frame after feature correlation (c.f. Sec. 4.3.3). This can be an auxiliary feature for the scene flow decoding. Although simple, this strategy, to our surprise, gives a non-trivial contribution to our performance with a decrease of 0.01m on overall EPE3D. Lastly, the biggest improvement (i.e., a 0.012m decrease on EPE3D) is brought by our constrained regression, in which the scene flow component on each axis is clamped by a fixed threshold (e.g. 0.1m). This is reasonable as non-viable results, for example, a scene flow vector with a length of 0.5m, can seriously degrade our results after averaging over all points.
6.3. Sensitivity Analysis
Here we investigate the sensitivity of our network against a) global aggregation mechanisms, b) test environments and c) the number of training subjects, and conduct separate analyses for them in the following.
EPE3D (m) $\downarrow$ Acc3D $\uparrow$
Mechanism  All  Moving  Static Strict Relax
Attention-based 0.046 0.051 0.009 0.406 0.703
Max-pooling 0.050 0.059 0.011 0.394 0.679
Average-pooling 0.047 0.055 0.009 0.399 0.698
Table 3. Comparison of global aggregation mechanisms.
Impact of global aggregation mechanism. Besides the attention-based mechanism, another two typical operations, i.e., max-pooling and average-pooling, are tested into our network for global feature aggregation (c.f. Sec. 4.3.2). The comparison results are shown in Table 3. From the table, the attention-based mechanism yields the best performance as it can learn to dynamically adjust the weights according to per-point local features. The average-pooling performs better than the max-pooling. We credit this to the ability of the average-pooling to retain the mean feature across all points, which is more robust to outliers.
EPE3D (m) $\downarrow$ Acc3D $\uparrow$
Scene  All  Moving  Static Strict Relax
Hallway 0.040 0.046 0.007 0.464 0.746
Square 0.042 0.049 0.009 0.435 0.722
Parking lot 0.051 0.058 0.011 0.374 0.671
Table 4. Results for different test environments.
Impact of test environments. In Table 4, we report the results of our trained model on each scene individually. As we can see, our model shows the best performance for the indoor scene (i.e., hallway) while performing worst in the parking lot scene. We credit the difference in performance between test environments to two factors. First, the background clutter in outdoor scenes, especially the parking lot (c.f. Fig. 4), is more complicated, which results in severer multi-path reflection during measurement. This decreases the SNR in radar data and further degrades the fidelity of generated radar point clouds. Second, our pseudo label generation (c.f. Sec. 4.4) is affected by the changeable outdoor illumination conditions. Not only does the 2D keypoint estimation on RGB images become more challenging, but also the depth measurements can be interfered with by the strong sunlight. Consequently, the quality of skeleton estimation in outdoor scenes is lower and the training is not as effective as in the indoor scene.
Figure 6. Impact of the number of training subjects.
Impact of training subjects number. In the above experiments, the data from 6 subjects are used for training (c.f. Sec. 5.1.2). Here we utilize 2/3/4/5 people out of the training set respectively for training to analyze the impact of the number of training subjects. As seen in Figure 6, the performance of our network increases in all metrics as more subjects are added for training. This proves that there is still room to improve our performance when more training data is available. We can also observe that, even with only 3 training subjects (9k data frames), our network can still generalize well to the testing set with an overall EPE3D of 0.066m and a relax Acc3D of 0.56. This result further demonstrates the feasibility of our automatic scene flow labelling scheme (c.f. Sec. 4.4) to generate reliable pseudo labels.

### 13 Mar 2024 v5

@@ 3.3 Technical Challenges
Figure 2: mmWave-based scene flow network architecture. The network takes consecutive radar point clouds as the input and outputs the scene flow in between.
3.4 Scene Flow Network
We employ end-to-end trainable deep neural networks for learning point-based scene flow estimation, in line with the state-of-the-art [22, 81, 43, 28]. The network architecture is sketched in Fig. 2. Particularly, we overcome the sparsity and noise challenges by integrating global features with local ones to provide a comprehensive view of each point cloud. Additionally, we address the lack of Doppler velocity and capture inconsistencies by leveraging a GRU network [18] to incorporate temporal information into the scene flow estimation. In subsequent sections, we will describe the details of each module in the network.
Local Feature Abstraction. Given mmWave radar point clouds $\mathcal{X}$ and $\mathcal{Y}$ as inputs, we encode their local features using four parallel SA layers [62] with varying grouping radii, enabling multi-scale local feature extraction to account for radar point clouds’ non-uniform density. Each SA layer generates a local feature $l_{i,s}$ for a given scale $s$, which is then transformed into a higher-level representation $k_{i,s}$ using a shared-weight MLP with scale-specific parameters. For both point clouds, we concatenate these high-level features across scales to form multi-scale local feature sets $\mathcal{K}^{\mathcal{X}}$ and $\mathcal{K}^{\mathcal{Y}}$ (cf. Fig. 2). Additionally, a separate context extractor, distinct from but structurally similar to the local encoder, generates context features $\mathcal{K}^{\mathcal{X}}_{c}$ for $\mathcal{X}$, enhancing the feature representation.
Global Feature Aggregation. In this module, the local feature vector of each radar point is first transformed into a scalar attention weight $w_{i}$ using an MLP. These weights are normalized to sum to 1 for numerical stability. The global feature vector $g$ is then derived by a weighted sum of all local features, offering an improvement over max-pooling by dynamically adjusting point-wise weights. This global vector is concatenated with each local feature to form local-global representations $\mathcal{Z}^{\mathcal{X}},\mathcal{Z}^{\mathcal{Y}}$ for $\mathcal{X}$ and $\mathcal{Y}$. Additionally, a global feature for the context $\mathcal{L}_{C}^{\mathcal{X}}$ is obtained through another MLP, resulting in $\mathcal{Z}^{\mathcal{X}}_{C}$, as seen in Fig. 2.
3.4.1 Flow Embedding Generation.
@@ 3.4 Scene Flow Network
Given the local-global features of two radar point clouds, i.e., $\mathcal{Z}^{\mathcal{X}},\mathcal{Z}^{\mathcal{Y}}$, we use the cost volume layer [83] to compute the correlation between them, as seen in Fig. 2. By aggregating the spatial relationship and feature similarities between two frames, the point motions are encoded into the cost volumes, denoted as $\mathcal{O}=\{o_{i}\}_{i=1}^{N}$. Thanks to the global feature aggregation, the holistic frame information can also be correlated, which yields more robust and stable costs. To further mix it with the context, we then stack the cost volumes $\mathcal{O}$, context features $\mathcal{L}_{C}^{\mathcal{X}}$ and pass the features into another local encoder to obtain the flow embedding =$\mathcal{B}=\{b_{i}\}_{i=1}^{N}$.
@@
Temporal Information Propagation. Directly propagating 2-dimension flow embedding from previous frames suffers from a) high computation overload, and b) the change in the number of points. To overcome these issues, we propose to temporally update its global vector $g^{\mathcal{B}}$ instead of the flow embedding itself $\mathcal{B}$ (cf. Fig. 2). We apply a GRU network [18] to update it as hidden states temporally and obtain the final global representation $h^{t}=\mathrm{GRU}(h^{t-1},g^{\mathcal{B}};\theta_{g})$, where $h^{t-1}$ is the global representation from the last frame and $\theta_{g}$ is the GRU network parameters. Lastly, we concatenate this updated global representation to each point in the flow embedding and denote the final features as $\mathcal{A}=\{a_{i}\}_{i=1}^{N}$.
@@
Constrained Scene Flow Regression. Given the final features $\mathcal{A}$ produced above, we can use an MLP-based flow regressor to decode per-point scene flow =$\mathcal{{F}}=\{{f}_{i}\}_{i=1}^{N}$. However, estimating unconstrained scene flow may lead to non-viable results, e.g., the magnitude of the flow vector exceeds the normal scale of human body movement. To constrain our predictions, we propose to clamp the estimated scene flow before returning it. We set a fixed threshold $\epsilon$ and constrain the scene flow on each axis to be within the range of [$-\epsilon,\epsilon$], as shown in Fig. 2.
3.5 Automatic Scene Flow Labelling
To address the human sensing scenarios where non-rigid motion dominates, we propose a cross-modal automatic scene flow labelling scheme (cf. Fig. 3), where pseudo scene flow labels are obtained from the 3D human skeletons. Our motivation is based on the observation that the non-rigid human body can be roughly segmented into multiple skeletons each of which can be seen as a rigid body, such as the neck and thigh bone. To simplify the problem of modelling human body motion, we attribute the human dynamics to the rigid motion of these skeletons and further assume the scene flow of points in the vicinity is induced by them.
@@ 3.5 Automatic Scene Flow Labelling
Human Skeleton Annotation. In human sensing applications, the human skeleton is usually characterised by the 3D position of its two endpoints, aka. keypoints. To quickly and conveniently annotate such keypoints, we refer to the RGB-D images recorded by the co-located RGB-D camera in this work. Specifically, we first utilize an open-source pose estimation library (e.g. OpenPose [14]) to label 2D keypoints on RGB images and then uplift each 2D keypoint to 3D using its corresponding depth value. Then the human skeleton labels can be obtained using intrinsically connected 3D keypoints, as shown in Fig. 3.
@@
Keypoint-Based Label Filtering. To reduce noise from the inaccuracies of 2D keypoint estimation, we filter pseudo scene flow labels using confidence scores and 3D displacement of keypoints. Keypoints with confidence below 0.5 are first removed, and those with displacement over 0.5m between frames are further filtered out. This process yields a set of valid keypoints per frame, and skeletons with both two end keypoints valid are considered valid. We then generate a mask $\mathcal{M}$, used below to minimise the impact of noise.
Figure 3: Automatic scene flow labelling pipeline. With the help of the co-located RGB-D camera, we first label 3D human skeletons and then generate noisy pseudo scene flow labels with respect to the skeleton-based rigid-motion assumption.
3.6 Loss Function
@@
$\mathcal{L}=\alpha_{l}\mathcal{L}_{large}+\alpha_{s}\mathcal{L}_{small}$ (1)
where $\alpha_{l}$ and $\alpha_{s}$ are hyperparameters that balance large- (i.e., larger than a threshold $\zeta$) and small-scale movement impacts, with the $L_{2}$ distance measuring prediction errors. This approach aims to prevent the network from settling into local minima by reducing the influence of static or minor-moving points.
4 Experiment
4.1 Dataset Collection
To facilitate the evaluation of our approach, we collect a large-scale multi-modal human motion sensing dataset with annotation labels for various tasks.
Platform. As exhibited in Fig. 4 (a), we use a commercial Vayyar vTrigB imaging mmWave radar [4] and a RealSense D455 depth camera [1] to capture mmWave radar point clouds and RGB-D images respectively. Both sensors are fixed on a collection board mounted on a tripod to ensure their relative position is unchanged once calibrated. We alternately query the data frame from two sensors and store them in a PC, resulting in synchronized multi-modal data. Note that the Vayyar radar we used is bespoken designed for fine-grained human sensing. It provides radar data with a higher resolution than most other mmWave radars, e.g., TI radars [70], thus enabling us to estimate fine-grained scene flow for radar points. Please find more details on this type of mmWave radar in the supplementary materials.
@@ 4.1 Dataset Collection
Procedure. 12 participants of various genders, ages and heights are recruited for data collection in this work [^1] The study has received the ethical approval from XXX (no name for blind review), and participant consent forms were signed before the collection.. We choose three sites as our human sensing experiment scenes, as shown in Fig. 4 (b). Each participant is asked to perform 5 ‘in-set’ activities for each scene, and 3 ‘out-of-set’ activities in one scene for the evaluation of the generalization to unseen activities (cf. Fig. 4 (d)). Every subject is asked to wear a face mask to protect their identities from being recognised. The distance between subjects and sensors is 2-4m and randomly selected by the participant during collection.
Statistics. After data collection, we crop each sequence to 200 frames for uniform activity distribution. The whole ‘in-set’ dataset consists of $12\times 3\times 5\times 200=$ 36k frames in total. To assess the generalization ability to new subjects, we divide the dataset into three parts by subject following the ratio of train:val:test = 3:1:2. The ’out-set’ testing set is composed of $12\times 3\times 200=$ 7.2k frames.
@@
Figure 4: Collection setup, test environment, subject activities and pseudo pose labels.
4.2 Evaluation Setup
Data Labelling. Our automatic labelling scheme (cf. Sec. 3.5) is used to annotate scene flow for the training and validation set by retaining 14 keypoints from OpenPose [14], leading to 13 connections between keypoints (cf. Fig. 4 (c)). Invalid skeletons are filtered out, and pseudo scene flow labels are assigned to points on valid skeletons. For the testing set, the same pipeline is used, but all points are annotated, with manual inspection and correction to 2D keypoints. For human activity recognition, activities are manually labelled during recording. Human parsing labels are derived from point-skeleton affiliations in scene flow labelling (cf. Fig. 3), with the valid mask $\mathcal{M}$ applied to filter invalid points during training. The human body part tracking labels are directly obtained from the 3D keypoints identified in the scene flow labelling process.
Evaluation Metric. For scene flow evaluation, we use the EPE3D (m) and Acc3D metrics following [10, 23]. Specifically, we redefined the Acc3D strict and relax requirements by reducing 0.05/0.1m to 0.025/0.05m to adapt to our human-centric scenario. The overall accuracy (OA) (%) is reported for both the human activity recognition and human parsing tasks, while mIoU (%) is used for HP only. To evaluate the human body part tracking, we compute the mean joint localization error (mJE) (m).
Table 1: Comparison of scene flow results between ours and state of the arts. $\uparrow$ means bigger values are better while $\downarrow$ means smaller values are better.
Table 2: Breakdown results of our scene flow network.
Table 3: Generalization of our model to new activities.
@@ 4.2 Evaluation Setup
FlowNet3D [52] 0.293 0.290 0.259 0.016 0.095
PPWC-Net [83] 0.171 0.181 0.128 0.138 0.179
Graph Prior [59] 0.315 0.322 0.283 0.007 0.011
FLOT [60] 0.299 0.307 0.265 0.015 0.094
FlowStep3D [43] 0.243 0.251 0.216 0.062 0.109
PV-RAFT [81] 0.161 0.170 0.107 0.179 0.292
RaFlow [23] 0.107 0.115 0.094 0.271 0.427
Bi-PFNet [17] 0.159 0.168 0.111 0.153 0.264